OpenStudy (goldphenoix):

(1/3)*(2/4)*(3/5)*(4/6)*( 5/7)*(6/8)*(7/9)*(8/10) = (1/10) * ? A: (3/19) B: (2/9) C: (1/9) D: (2/90)

OpenStudy (whpalmer4):

\[\frac{1*2*3*4*5*6*7*8}{3*4*5*6*7*8*9*10} = \frac{1}{10}*x =\frac{1*2*\cancel{3*4*5*6*7*8}}{\cancel{3*4*5*6*7*8}*9*10} \]

OpenStudy (goldphenoix):

If only I was smart as you. I don't think outside the box. T~T So let see, the numerator is 2. Denominator is 90 So it's (2/90)?

OpenStudy (goldphenoix):

Oh wait. I chose D, but I got it wrong.

OpenStudy (loser66):

bow whpalmer4 can I be your student?

OpenStudy (whpalmer4):

No, after cancellation we have \[\frac{1*2}{9*10} = \frac{1}{10}*x\]

OpenStudy (goldphenoix):

Oh I see. \[\large \large \frac{ 2 }{ 90 } = \frac{ 1 }{ 10 }x\] Let see if I can remember this. We times both side by 10?

OpenStudy (whpalmer4):

Wait! Stop! Look! \[\frac{1*2}{9*10} = \frac{1}{10}*x\] \[\frac{1*2}{10*9} = \frac{1}{10}*x\]

OpenStudy (goldphenoix):

I don't get it. Isn't it the same thing?

OpenStudy (whpalmer4):

Prepare your forehead for imminent impact :-) \[\frac{1*2}{10*9} = \frac{1}{10}*x\]\[\frac{1}{10}*\frac{2}{9} = \frac{1}{10}*x\]

OpenStudy (goldphenoix):

o.O?

OpenStudy (whpalmer4):

\[\cancel{\frac{1}{10}*}\frac{2}{9} = \cancel{\frac{1}{10}}*x\]

OpenStudy (goldphenoix):

OH WOW! T~T

OpenStudy (goldphenoix):

I wasn't prepared for that...

OpenStudy (whpalmer4):

Don't worry, those forehead dents go away quickly :-)

OpenStudy (goldphenoix):

I hope so. T~T Thanks! :D

OpenStudy (whpalmer4):

you're welcome!